OpenStudy (anonymous):

Solve and Check. 2/3V=6 A 1/9 B 1/4 C 9 D 4

myininaya (myininaya):

is that \[\frac{2}{3}v=6\]

myininaya (myininaya):

If so try multiplying both sides by the reciprocal of 2/3

myininaya (myininaya):

That is the usual way of solving these types. When you have \[ \\ \text{ assume } a \neq 0 \text{ and } b \neq 0 \\ \frac{a}{b}x=c \\ \text{ you multiply both sides by } \frac{b}{a} \\ \frac{\cancel{b}}{\cancel{a}} \frac{\cancel{a}}{\cancel{b}}c=\frac{b}{a}c \\ \text{ so } x=\frac{b}{a}c\]

OpenStudy (anonymous):

I up dated it so there

myininaya (myininaya):

That still doesn't change the process You still need to multiply both sides by the reciprocal of 2/3

OpenStudy (anonymous):

ok thanks but how would i find my answer

myininaya (myininaya):

Multiplying both sides by the reciprocal of 2/3

myininaya (myininaya):

try doing that as a first step

OpenStudy (anonymous):

ok so i got 1 by mutiplying 2/3*3/2

myininaya (myininaya):

right so 2/3*3/2*v=1*v=v but you also need to multiply the other side by 3/2

OpenStudy (anonymous):

so v is gonna = 4 or 9 right

myininaya (myininaya):

\[\frac{2}{3}v=6 \\ \\ \text{ multiply both sides by } \frac{3}{2} \\ \frac{3}{2} \cdot \frac{2}{3} v=6 \cdot \frac{3}{2} \\ 1v=6 \cdot \frac{3}{2} \\ v=\frac{6 \cdot 3}{2}\] it will only be one answer

OpenStudy (anonymous):

So the answer is 1/9

OpenStudy (anonymous):

or would it be 9

myininaya (myininaya):

what do you think 6*3 is ?

myininaya (myininaya):

what do you think 18 divided by 2 is ?